# A Self-Supervised Model for Multi-modal Stroke Risk Prediction

### v1

@@ Abstract
Predicting stroke risk is a complex challenge that can be enhanced by integrating diverse clinically available data modalities. This study introduces a self-supervised multimodal framework that combines 3D brain imaging, clinical data, and image-derived features to improve stroke risk prediction prior to onset. By leveraging large unannotated clinical datasets, the framework captures complementary and synergistic information across image and tabular data modalities. Our approach is based on a contrastive learning framework that couples contrastive language-image pretraining with an image-tabular matching module, to better align multimodal data representations in a shared latent space. The model is trained on the UK Biobank, which includes structural brain MRI and clinical data. We benchmark its performance against state-of-the-art unimodal and multimodal methods using tabular, image, and image-tabular combinations under diverse frozen and trainable model settings. The proposed model outperformed self-supervised tabular (image) methods by 2.6% (2.6%) in ROC-AUC and by 3.3% (5.6%) in balanced accuracy. Additionally, it showed a 7.6% increase in balanced accuracy compared to the best multimodal supervised model. Through interpretable tools, our approach demonstrated better integration of tabular and image data, providing richer and more aligned embeddings. Gradient-weighted Class Activation Mapping heatmaps further revealed activated brain regions commonly associated in the literature with brain aging, stroke risk, and clinical outcomes. This robust self-supervised multimodal framework surpasses state-of-the-art methods for stroke risk prediction and offers a strong foundation for future studies integrating diverse data modalities to advance clinical predictive modeling.
@@ 1 Introduction
Stroke ranks as the second leading cause of death worldwide, responsible for 11.6% of global fatalities in 2019. It often results in neurological damage and long-term disability in adults, imposing significant health and economic challenges Pu et al. [2023], Feigin et al. [2021]. Early detection through predictive models is crucial in preventing severe outcomes, as cerebrovascular events can cause irreversible brain damage within hours Flora and Nayak [2019]. The complexity of stroke, driven by multiple risk factors, highlights the importance of integrating multi-modal data to improve diagnostic accuracy and treatment strategies. Among the various imaging techniques, Magnetic Resonance Imaging (MRI) stands out as a highly effective tool, offering high-resolution, non-invasive assessments of structural abnormalities and detailed visualization of the brain’s vascular network Hartwig et al. [2009].
@@
Uni-modal predictive models Prior works mainly use convolutional neural networks (CNN) that can leverage the high-dimensional imaging information for diagnosing patients Zhang and Qie [2023]. Yu et al. applied deep learning algorithms to extract meaningful imaging features in an increasing order of hierarchical complexity to make predictions of the infarct volume Yu et al. [2020]. Other models that use only clinical data, often assume linear relationships between traditional risk factors such as age, gender, smoking status, blood pressure, diabetes, cholesterol levels, and body mass index Hippisley-Cox et al. [2024], An et al. [2020], You et al. [2023]. Alaa et al. used AutoPrognosis, an ensemble machine learning approach, to outperform conventional models like the Framingham score and Cox models Alaa et al. [2019]. A major limitation of these models is that they don’t integrate complementary information from other modalities, similar to how clinicians diagnose using multiple data sources. Biobanks like the UK Biobank (UKB) have become invaluable in this context, providing vast datasets integrating imaging and clinical information to train machine learning models for disease prediction Littlejohns et al. [2020], UK BioBank [2021].
@@
Multi-modal predictive models Several studies have employed multi-modal data to improve diagnostic capabilities by integrating diverse data types Liu et al. [2015]. For example, MultiSurv model has shown success by fusing image and tabular data for cancer survival prediction Vale-Silva and Rohr [2021]. In another study, integration of retinal images and clinical data was leveraged to improve cardiovascular disease prediction Huang et al. [2024]. Multi-modal models combining image and clinical data have demonstrated better prediction performance for disability prediction in stroke patients White et al. [2023], Liu et al. [2023]. However, CNNs tend to prioritize image features, and simple image-tabular CNN concatenation fails to enhance predictive models due to insufficient cross-modal interactions. To address this, Wolf et al. developed the Dynamic Affine Feature Map Transform (DAFT), which conditions convolutional feature maps on both image and tabular data, enabling a two-way information exchange via an auxiliary neural network Wolf et al. [2022]. While DAFT reduces issues related to the large number of trainable parameters in standard 3D CNNs and the curse of dimensionality, it may sacrifice some predictive power compared to deeper models like ResNet. Although recent models show promise in biomedical prediction tasks, their clinical translation is hindered by limited annotated datasets, low disease prevalence, and the risk of overfitting. Self-supervised learning (SSL) is a powerful technique for extracting representative features from unlabeled data, making it valuable for early disease risk identification.
@@
Self-supervised models Unlike traditional supervised learning, SSL defines pretext tasks that allow models to learn meaningful representations from raw data Balestriero et al. [2023]. One prominent SSL technique is contrastive learning, which trains encoders to generate augmented views of a sample, maximizing similarity between these views while minimizing similarity with other samples Balestriero et al. [2023]. Popular methods such as SimCLR Chen et al. [2020], BYOL Grill et al. [2020], and MOCO He et al. [2020] have demonstrated success in imaging tasks, while VIME Houthooft et al. [2016] and SCARF Bahri et al. [2022] are leading approaches for tabular data. Emerging approaches, like contrastive language-image pre-training (CLIP) strategy, have evolved from unimodal methods to integrate diverse modalities. While there was an extensive work done for cardiovascular diseases prediction Radhakrishnan et al. [2023], Du et al. [2024], Hager et al. [2023], Girlanda et al. [2024], stroke risk prediction through volumetric brain images and clinical health records remains underexplored.
@@
We present for the first time, to the best of our knowledge, a self-supervised multi-modal approach integrating 3D brain MRIs with clinical tabular data for stroke risk prediction. As depicted in Figure 1, our methodology incorporates cross-modal interactions via CLIP loss Radford et al. [2021] and image-tabular matching (ITM) loss Li et al. [2021], Du et al. [2024]. We demonstrate that our learning strategy outperforms leading (self-)supervised unimodal methods and that multi-modal image-tabular pre-training leads to better representations and improved downstream performance. Lastly, we validate the model’s learned features through visual activation maps, which align with established clinical and neurological findings on stroke-related brain pathology. Code is available at https://github.com/CamilleDelgrange/SSMSRPM.
@@ 2.1 Dataset
Our analyses are performed on T2-Fluid Attenuation Inversion Recovery (FLAIR) brain volumes, and over a subset of clinical information spanning across five categories, extracted from the UKB: demographics, lifestyle, biomarkers, comorbidities, and medication. The complete list of features is available in the supplementary materials. Continuous features are standardized using z-score normalization, while categorical data is one-hot encoded. For our experiments, we use 5000 and 500 samples for training and validation sets respectively for the model pre-training stage. Train, validation, and test subset for the downstream fine-tuning stage use 278, 93, and 93 samples respectively. The fine-tuning sets are stratified according to sex, age and stroke diagnosis to account for confounders to avoid spurious correlations and class imbalance. To handle missing tabular data, we use an iterative multivariate imputer based on Multivariate Imputation by Chained Equations (MICE) van Buuren and Groothuis-Oudshoorn [2011], modelling missing features as a function of existing features over multiple imputation rounds. Missing categorical data is replaced by the most frequent category. This step is performed after data normalization, to ensure that the means and standard deviations are calculated only from recorded values. The 3D brain images are registered to Montreal Neurological Institute brain template (MNI) space, have uniform dimensions of $182\times 218\times 182$ and a voxel size of $1mm^{3}$ and are processed using the UKB imaging pipeline Alfaro-Almagro et al. [2018b]. Key image-derived phenotypes (IDPs), such as segmented brain tissue volumes and white matter hyperintensity (WMH) volumes, are extracted and used as brain IDPs. Brain lesion segmentation is performed using the BIANCA tool to produce 3D binary lesion masks Alfaro-Almagro et al. [2018a]. Furthermore, lesion segmentation masks are characterized by pyradiomics Van Griethuysen et al. [2017] through radiomic features such as volume, area, elongation, and sphericity and these features are used as lesion IDPs.
@@ 2.2 Multi-modal self-supervised framework
Our pipeline is split into two sequential steps. First, we pre-train the tabular and imaging encoders (Figure 1 A) and then we fine-tune them with labels from downstream task (Figure 1 B). Each batch of data contains pairs of imaging $x_{j_{i}}$ and tabular $x_{j_{t}}$ samples. These samples are augmented by random transformations $t\sim\tau$ from a set of parametric transforms $\tau$, such as random cropping and affine transforms for the images, or random feature corruption for the tabular data. We use an image augmentation rate of 95% for the model to still occasionally see unaltered data to capture the original data distribution for transfering the learnt features to the downstream task. The corruption rate of the tabular data is set to 0.3 as in the original tabular method SCARF Bahri et al. [2022]. For a given reference point, known as anchor $x$, the positive samples are the ones derived from $x$ transformations while other samples in the batch are considered as negative samples. Augmented images $x_{j_{i}}$ and tabular data $x_{j_{t}}$ are passed through the imaging encoder $f_{\theta_{I}}$ and tabular encoder $f_{\theta_{T}}$ to generate the embeddings. These embeddings are propagated through the separate projection heads $f_{\phi_{I}}$ and $f_{\phi_{T}}$, and brought into a shared latent space as projections $z_{j_{i}}$ and $z_{j_{t}}$, which are L2-normalized onto a unit hypersphere. The projections are pulled and pushed in the shared latent space according to the CLIP loss Radford et al. [2021], which maximizes the cosine similarity of projections from the positive samples and minimizes the similarity of projections from the negative samples in the batch. In contrast to the original InfoNCE loss used in SimCLR Chen et al. [2020], and following the CLIP loss, the projected embeddings similarities are contrasted between data modalities. An image projection is therefore defined as :
@@
Considering all N subjects in a batch, the loss for the imaging modality is defined as follows:
@@
where $\tau$ is the temperature parameter. In our experiments, a temperature of 0.1 is selected to work best, following Chen et al. [2020]. $l_{t,i}$ is computed analogously and CLIP loss is defined as follows:
@@
We choose value of 0.5 for the $\lambda$ as regularization parameter. The aim is to learn patient-wise representations invariant to the variation of the image-tabular pairs. Hard negative samples are crucial in contrastive learning as they help the model distinguish between similar samples, preventing trivial solutions and enhancing its robustness. We implement a hard negative mining strategy to predict whether image-tabular data pairs are positive or negative, using image-tabular matching (ITM) loss. In this approach, for each image or tabular representation, we identify an unmatched tabular or image representation from the mini-batch. This selection is based on similarity scores computed using the CLIP method, which serves as the sampling weight for the negative pairs Du et al. [2024], Li et al. [2021]. A multi-modal interaction module is introduced, as shown in Figure 1, which takes the output of the projector heads to perform inter-modality learning and generates a multi-modal representation. It uses a cross-attention mechanism Vaswani et al. [2017], enabling tabular embeddings to attend to relevant image embeddings. The multi-modal interaction module contains two transformer layers, with four attention heads and a hidden dimension of 256, each including self-attention, cross-modal attention, an MLP feed-forward module and layer normalization Du et al. [2024]. The output of the multi-modal module is a [CLS] token, aggregating the information from the entire sequence, used for downstream classification task, where the model needs a single feature vector representing the entire input Ye et al. [2023]. The [CLS] embedding is capturing a joint representation of the image-tabular pair that is fed into the ITM predictor (a linear layer) to match the prediction based on a binary cross-entropy loss $\mathcal{L}_{ITM}$. Therefore, the complete loss is expressed as:
@@
Downstream task predictions After pre-training, the projection heads were replaced by fully connected layers. Extracting the representation before the projector has been shown to improve downstream tasks performance Balestriero et al. [2023]. For downstream fine-tuning and binary classification of healthy versus stroke (Figure 1 B), we employ ensemble learning to improve model generalization and performance by leveraging the rich representations from the image encoder, tabular encoder, and the multi-modal transformer interaction module. All pre-trained models are evaluated using linear probing (frozen) and fine-tuning (trainable). The frozen models use tuned linear classifiers after the feature extractors. The datasets used for model fine-tuning are balanced in each batch of training, validation, and test subset. This way, we reduce potential bias due to class-imbalance, as well as unstable and slow training due to imbalance batch distributions.
Figure 1: Pipeline for joint imaging and tabular data pre-training (A) and downstream fine tuning (B). CLIP loss is applied on projected data to align the image and tabular representations. Hard negative pairs are mined through CLIP similarities within the batch. A transformer block with self-attention and cross-attention layers is used to cross-attend both modalities, resulting in a multi-modal [CLS] token fed to a classifier and used for further downstream fine tuning. Image-Tabular Matching (ITM) loss evaluates the image-tabular pair matching. In the downstream task, an ensemble classifier is fine-tuned to predict healthy versus stroke from pre-trained imaging, tabular and multimodal encoders.

### v2

# bibliography.bib

@@ Abstract
Predicting stroke risk is a complex challenge that can be facilitated by integrating diverse clinically available data modalities. This study introduces a self-supervised multi-modal framework combining brain MRI and clinical records to improve stroke risk prediction prior to onset. By leveraging large unannotated clinical datasets, the framework captures complementary and synergistic information across image (brain MRI) and tabular (clinical records) data modalities, serving as a multi-modal foundation model. Our approach employs a contrastive learning framework integrating language-image pretraining with an image-tabular matching module, to better capture the multimodal characteristics of each patient. We benchmark its performance against state-of-the-art uni-modal and multi-modal methods across tabular, image and combined data modalities under various model settings. The proposed multi-modal foundation model outperformed self-supervised tabular (image) methods by 2.1% (2.8%) in ROC-AUC and by 10.6% (12.8%) in balanced accuracy. Additionally, it showed a 7.6% increase in accuracy compared to the best multi-modal supervised model. Using interpretable tools, our approach demonstrated better integration of tabular and image data, providing richer and more aligned embeddings. Moreover, interpretable model-derived heatmaps further underscored brain regions frequently associated in the literature with brain aging, stroke risk, and clinical outcomes. This robust self-supervised multi-modal foundation model surpasses state-of-the-art methods for stroke risk prediction and lays the groundwork for future studies integrating diverse data modalities to advance clinical predictive modeling. Our approach is the first to allow the prediction of stroke risk prior to onset by leveraging multimodal clinical data, and could support earlier identification of at-risk patients and facilitate more proactive clinical management.
@@ 1 Introduction
Stroke ranks as the second leading cause of death worldwide, responsible for 11.6% of all deaths globally in 2019 [Pu2023, Feigin2021]. It often results in neurological damage and long-term disability in adults, imposing significant health and economic challenges [Pu2023, Feigin2021]. Early identification of patients at highest risk for stroke is crucial to prevent such events and long-term poor outcomes [Flora2019]. The complexity of stroke, driven by multiple risk factors, highlights the importance of integrating multi-modal data to improve prediction accuracy and anticipate treatment strategies [durso2024probabilistic]. Among the various imaging modalities, Magnetic Resonance Imaging (MRI) is uniquely suited to identify brain biomarkers potentially indicative of future stroke events. With its high-resolution, non-invasive capabilities, MRI enables precise assessment of structural abnormalities and detailed visualization of the brain’s vascular network, offering critical insights into underlying conditions that may contribute to stroke risk [Hartwig2009].
@@
Uni-modal predictive models Predictive models play a crucial role in healthcare, as early identification of patients at risk is essential to prevent irreversible brain and body damage and ensure that they receive optimal care. Initially, these models were developed using a single clinical modality (uni-modal) such as images or clinical information from patient health records. Previous works mainly use convolutional neural networks (CNN) that can leverage high-dimensional imaging information for diagnosing patients [Zhang2023]. Yu et al. applied deep learning algorithms to extract meaningful imaging features in an increasing order of hierarchical complexity to make predictions of the infarct volume [yu2020use]. Other models that use only clinical records, often assume linear relationships between traditional risk factors such as age, gender, smoking status, blood pressure, diabetes, cholesterol levels, and body mass index [Hippisley-Cox2024, An2020, You2023]. Alaa et al. used AutoPrognosis, an ensemble machine learning approach, to outperform conventional models like the Framingham score and Cox models [Alaa2019]. A major limitation of these models is their inability to integrate complementary information from other modalities, unlike clinicians who rely on multiple data sources for diagnosis. Biobanks like the UK Biobank (UKB) has become invaluable in this context, providing vast datasets integrating imaging and clinical information to train machine learning models for disease prediction [Littlejohns2020, UKBioBank2021].
@@
Multi-modal predictive models Several studies have integrated diverse data types to improve diagnostic capabilities [Liu2015]. For example, de Havenon et al. [deHav2023improvement] highlighted the value of imaging biomarkers by improving cerebrovascular events prediction through the incorporation of white matter hyperintensities derived from brain imaging into clinical records, emphasizing the potential of such approaches to refine risk assessments. MultiSurv has shown success by fusing image and tabular data for cancer survival prediction [Vale-Silva2021]. In another study, integration of retinal images and clinical records was leveraged to improve cardiovascular disease prediction [huang2024predictingstrokeretinalgraphs]. Multi-modal models combining imaging data and clinical records have demonstrated better prediction performance for disability prediction in stroke patients [White2024, Liu2023]. However, CNNs tend to prioritize image features, and simple image-tabular CNN concatenation fails to enhance predictive models due to insufficient cross-modal interactions. To address this, Wolf et al. developed the Dynamic Affine Feature Map Transform (DAFT), which conditions convolutional feature maps on both image and tabular data, enabling a two-way information exchange via an auxiliary neural network [Wolf2022]. While DAFT reduces issues related to the large number of trainable parameters in standard 3D CNNs and the curse of dimensionality, it may sacrifice some predictive power compared to deeper models like ResNet. Although recent models show promise in biomedical prediction tasks, their clinical translation is hindered by limited annotated datasets, low disease prevalence, and the risk of overfitting. Self-supervised learning (SSL) is a powerful technique for extracting representative features from unlabeled data, making it valuable for early disease risk identification.
@@
Self-supervised models Unlike traditional supervised learning, SSL defines pretext tasks allowing models to learn meaningful representations from raw data [Balestriero2023]. One prominent SSL technique is contrastive learning, which trains encoders to generate augmented views of a sample, maximizing similarity between these views while minimizing similarity with other samples [Balestriero2023]. Popular methods such as SimCLR [Chen2020], BYOL [Grill2020], and MOCO [He2020] have demonstrated success in imaging tasks, while VIME [Houthooft2016] and SCARF [Bahri2022] are leading approaches for tabular data. Emerging approaches, like contrastive language-image pre-training (CLIP) strategy, have evolved from uni-modal methods to integrate diverse modalities. While there was an extensive work done for cardiovascular diseases prediction [Radhakrishnan2023, Du2024, Hager2023, girlanda2024enhancing], stroke risk prediction through volumetric brain images and clinical health records remains underexplored.
@@
We present for the first time, to the best of our knowledge, a scalable self-supervised multi-modal approach integrating research quality T2-FLAIR MRIs pre-stroke, along clinical records for stroke risk prediction, prior to onset. As depicted in Figure 1, our methodology incorporates cross-modal interactions via CLIP loss [Radford2021] and image-tabular matching (ITM) loss [Li2021, Du2024]. We demonstrate that our learning strategy outperforms leading (self-)supervised uni-modal methods and that multi-modal image-tabular pre-training aligns data modalities in a shared space improving representations, as well as downstream task performance. Lastly, we validate the model learned features through visual activation maps, which align with established clinical and neurological findings on stroke-related brain pathology. Our model is practical for real-world clinical use because it leverages a small dataset without the need of expert annotations, and could facilitate early identification of at-risk patients for a more proactive clinical management. The code repository for this project is available at the following link: https://github.com/CamilleDelgrange/SSMSRPM/.
@@ 2.1 Dataset
2.1.1 UKB Dataset
Our analyses are performed on the UKB, using a single (pre-stroke) T2-Fluid Attenuation Inversion Recovery (FLAIR) scan with timewise corresponding clinical patient information spanning across five categories: demographics, lifestyle, biomarkers, comorbidities, and medication. T2-FLAIR scans serve as an effective MRI sequence for identifying white matter hyperintensities (WMHs), which are strongly associated with increased stroke risk and serve as critical markers of cerebrovascular damage that may precede and predict stroke occurrence [Shim2015, Debette2010, Liou2009, Leiva-Salinas2010]. Stroke patients were identified using International Classification of Disease (ICD10) codes specific to stroke [Woodfield2015, Category43UKB] (ICD- 9 codes 430, 431, 434 and 436; ICD- 10 codes I60, I61, I63 and I64). ICD10 codes are maintained by the World Health Organization, used to record diagnoses during hospital admissions, and made available through the UKB. The stroke onset time was defined as the time between the baseline first imaging visit and the stroke onset. For healthy participants, they are defined as still healthy (e.g. without any stroke event) at the (second) repeated imaging timepoint. A summary of the cohort demographics including time-dependent variables can be found in Table LABEL:tab:demographics. The time variables were not used for training or inference however they give insights into the mean stroke onset time (from scan acquisition until the recorded clinical stroke event) as well as the last timepoint where the patient attended an imaging visit as healthy. The complete list of features used is available in Tables S1 and S2 in the supplementary materials. Continuous features are standardized using z-score normalization, while categorical data are one-hot encoded. For our experiments, as indicated in Table 2 during the model pre-training stage, we use 5000 and 500 samples for the training and validation sets, respectively. For the downstream fine-tuning stage, the training, validation, and test subsets -none of which were used during the pre-training stage- include 278, 93, and 93 samples, respectively. The fine-tuning sets are stratified by sex, age, and stroke diagnosis to address class imbalance; the splits were designed to preserve the distribution of these variables as observed in the pre-training dataset. To handle missing tabular data, we use an iterative multivariate imputer based on Multivariate Imputation by Chained Equations (MICE) [VanBuuren2011], modelling missing features as a function of existing features over multiple imputation rounds. Missing categorical data is replaced by the most frequent category. This step is performed after data normalization, to ensure that the means and standard deviations are calculated only from recorded values.
2.1.2 Brain MRI pre-processing
The 3D brain MRI images are registered to the Montreal Neurological Institute (MNI) brain template space [MNI], standardized to dimensions of $182\times 218\times 182$ voxels with a voxel size of $1~{}\mathrm{mm}^{3}$, and processed using the UK Biobank imaging pipeline [Alfaro-Almagro2018]. Key image-derived phenotypes (IDPs), such as brain tissue and white matter hyperintensity (WMH) volumes are extracted and referred to as “brain IDPs.” Segmentation of pre-stroke brain lesion abnormalities is performed using the BIANCA tool, generating 3D binary lesion masks [Alfaro-Almagro2018]. Additionally, features such as volume, area, and elongation are extracted from the lesion segmentation masks using pyradiomics [van2017computational]; this set of features is referred to as “lesion IDPs” throughout the paper.
Table 1: Table 1. Characteristics of the inference cohort for healthy individuals and stroke patients.
Category Feature (unit) Stroke Healthy $p$-value
Demographics Sex (Male, %) 60.87% 51.06% 0.4572
Age (years) 67.70 ± 7.30 61.38 ± 9.68 0.0006
BMI ($kg/m^{2}$) 28.01 ± 5.00 25.79 ± 3.82 0.0211
Waist Circumference (cm) 93.34 ± 14.37 86.68 ± 13.14 0.0247
Weight (kg) 80.86 ± 16.76 74.94 ± 15.21 0.7310
Time Variables Stroke Onset (months) 26.20 ± 18.79 N/A
Healthy Window (months) N/A 29.93 ± 10.30
Table 2: Summary of dataset splits used during the pre-training and fine-tuning stages.
Stage Subset Number of Samples
Pre-training Training 5000
Validation 500
Fine-tuning Training 278
Validation 93
Test 93
@@ 2.2 Multi-modal self-supervised framework
Our pipeline is split into two sequential steps. First, we pre-train the tabular and imaging encoders (Figure 1 A) and then we fine-tune them with labels from the downstream task (Figure 1 B). The image encoder consists of a ResNet-50 backbone [He2020], which outputs feature embeddings of size 512. The tabular encoder is a multi-layer perceptron (MLP) composed of two fully connected layers, each with an embedding dimension of 512 neurons. Each layer is followed by batch normalization, ReLU activation, and dropout with a rate of 0.3. Kaiming weight initialization is applied, and pretrained weights can optionally be loaded and frozen. The MLP encodes structured tabular data into dense embeddings. Both encoders are followed by a projection head, as used in SimCLR [Chen2020], which maps the embeddings to a lower-dimensional space for contrastive learning (128). The projection head consists of a fully connected hidden layer with ReLU activation, followed by a second linear layer without activation. This two-layer MLP allows the model to learn more expressive representations in the contrastive space, decoupled from the original embedding space. Each batch of data contains pairs of imaging $x_{j_{i}}$ and tabular $x_{j_{t}}$ samples which are augmented by random transformations $t\sim\tau$ from a set of parametric transforms $\tau$, such as random cropping and affine transforms for the images, or random feature corruption for the tabular data. We use an image augmentation rate of 95% for the model to still occasionally see unaltered data to capture the original data distribution for transfering the learnt features to the downstream task. The corruption rate of the tabular data is set to 0.3 as in the original SCARF method [Bahri2022].
2.2.1 Model pre-training
For a given reference point, known as anchor $x$, the positive samples are the ones derived from $x$ with transformations while other samples in the batch are considered as negative samples. Augmented images $x_{j_{i}}$ and tabular data $x_{j_{t}}$ are passed through the imaging $f_{\theta_{I}}$ and tabular encoder $f_{\theta_{T}}$ to generate the embeddings. These embeddings are propagated through the separate projection heads $f_{\phi_{I}}$ and $f_{\phi_{T}}$, and brought into a shared latent space as projections $z_{j_{i}}$ and $z_{j_{t}}$, which are l2-normalized onto a unit hypersphere. The projections are pulled and pushed in the shared latent space according to the CLIP loss [Radford2021], which maximizes the cosine similarity of projections from the positive samples and minimizes the similarity of projections from the negative samples in the batch. In contrast to the original InfoNCE loss used in SimCLR [Chen2020], and following the CLIP loss, the projected embeddings similarities are contrasted between data modalities. An image projection is therefore defined as in Equation 1:
@@ 2.2.1 Model pre-training
Considering all N subjects in a batch, the loss for the imaging modality $l_{i,t}$ is defined as follows, in Equation 2:
@@
where $\tau$ is the temperature parameter. In our experiments, a temperature of 0.1 is selected to work best, following [Chen2020]. $l_{t,i}$ is computed analogously and the CLIP loss is defined in Equation 3 as follows:
@@
[…]
We choose a value of 0.5 for the regularization parameter $\lambda$. The goal is to learn patient-specific representations that remain consistent despite variations in the image-tabular pairs. Hard negative samples are crucial in contrastive learning as they help the model to distinguish between similar samples, preventing trivial solutions and enhancing its robustness. We implement a hard negative mining strategy to predict whether image-tabular data pairs are positive or negative, using the image-tabular matching (ITM) loss. In this approach, for each image or tabular representation, we identify an unmatched tabular or image representation from the mini-batch. This selection is based on similarity scores computed using the CLIP method, which serves as the sampling weight for the negative pairs [Du2024, Li2021]. A multi-modal interaction module is introduced, as shown in Figure 1, which takes the output of the projector heads to perform inter-modality learning and generates a multi-modal representation. It uses a cross-attention mechanism [Vaswani2017], enabling tabular embeddings to attend to relevant image embeddings. The multi-modal interaction module contains two transformer layers, with four attention heads and a hidden dimension of 256, each including self-attention, cross-modal attention, an MLP feed-forward module and layer normalization [Du2024]. The output of the multi-modal module is a classification token [CLS], typically placed at the start of the input sequence to represent aggregated contextual information, used for downstream classification task, where the model needs a single feature vector representing the entire input [Oh2017]. The [CLS] embedding is capturing a joint representation of the image-tabular pair that is fed into the ITM predictor (namely, a linear layer) to match the prediction based on a binary cross-entropy loss $\mathcal{L}_{ITM}$. Therefore, the complete loss $\mathcal{L}$ is expressed in Equation 4, as:
@@
2.2.2 Model fine-tuning: downstream task predictions
After pre-training, the projection heads are replaced by fully connected layers. Extracting the representation before the projector has been shown to improve downstream tasks performance [Balestriero2023]. For downstream fine-tuning and binary classification of samples into healthy versus stroke (Figure 1 B), we employ ensemble learning to improve model generalization and performance by leveraging the rich representations from the image encoder, tabular encoder, and the multi-modal transformer interaction module. All pre-trained models are evaluated using linear probing (frozen) and fine-tuning (trainable). The frozen models use tuned linear classifiers after the feature extractors. The datasets used for model fine-tuning are balanced in each batch of training, validation, and test subset. This way, we reduce potential bias due to class-imbalance, as well as unstable and slow training due to imbalance batch distributions. We report test set results in Table 4. However, we also report test set results reflecting the original (imbalanced) stroke distribution (0.52% in the original UKB pre-training dataset), using a bootstrapping with replacement approach. These results are presented in Table 5 and are representative of the distribution in the UKB dataset used in this study. Additionally, we provide 95% confidence intervals derived from this analysis.
Figure 1: Pipeline for joint imaging and tabular data pre-training (A) and supervised fine tuning (B). CLIP loss is applied on projected data to align the image and tabular representations. Hard negative pairs are mined through CLIP similarities within the batch. A transformer block with self-attention and cross-attention layers is used to cross-attend both modalities, resulting in a multi-modal [CLS] token fed to a classifier and used for further downstream fine tuning. Image-Tabular Matching (ITM) loss evaluates the image-tabular pair matching. In the downstream task, an ensemble classifier is fine-tuned to predict healthy versus stroke from pre-trained imaging, tabular and multi-modal encoders.